# Adapt Tool Schemas to the Models

### v1

@@ 4.1 Benchmarks
MetaTool (Huang et al., 2024) evaluates tool selection capabilities across 4,287 test cases in four subtasks: (1) Similar tests semantic comprehension by distinguishing tools with overlapping functionalities (e.g., Sudoku vs. Tic-Tac-Toe); (2) Scenario select appropriate tools based on user-specific contexts and requirements (e.g., software engineers, students); (3) Reliability assesses whether models can directly indicate when no suitable tool is available rather than hallucinating; and (4) Multi-tool measures whether models can correctly select multiple tools when tasks require composition of functionalities.
@@
RoTBench (Ye et al., 2024) evaluates whether models can robustly use tools when descriptions contain various noise perturbations (e.g. reversed names). However, we focus exclusively on the Clean environment to isolate the effects of PA-Tool on core tool-use capabilities rather than noise robustness. We evaluate on 175 test cases across 568 tools in both single-turn and multi-turn settings. RoTBench measures two fundamental tool use capabilities: (1) Tool Selection evaluates whether models correctly identify the appropriate tool; and (2) Parameter Identification assesses whether models accurately extract the required parameter set, conditioned on correct tool selection.
4.2 Models and Baselines
We focus on models with 8B parameters or fewer (SLMs), evaluating four open-source models: Qwen2.5-3B/7B (Qwen et al., 2025), Llama3.1-8B (Meta AI, 2024a), and Llama3.2-3B (Meta AI, 2024b). For each model, we evaluate three configurations: Base without any modifications, Greedy using greedy decoding for schema generation, and PA-Tool with our proposed method. To contextualize SLM performance and quantify PA-Tool’s improvements, we also compare against three state-of-the-art closed-source models: GPT-4.1-mini (OpenAI, 2024), Gemini 2.5 Flash (Basu Mallick et al., 2025), and Claude Sonnet 4.5 (Anthropic, 2025).
4.3 Implementation Details
When generating tool schemas with PA-Tool, we generate 32 candidates with temperature 0.4 and set $\alpha$ to 0.2. For all benchmark evaluations, we employ a one-shot prompting strategy where the prompt includes: (1) the task instruction, (2) the target tool description (not the complete schema), and (3) one demonstration example. We use greedy decoding to ensure reproducible results across all experiments. We use accuracy as the primary evaluation metric across both benchmarks, measuring the percentage of test cases where the model’s predictions exactly match the ground-truth labels. Detailed prompt templates are provided in Appendix B.
5 Main Results
We present our main experimental results demonstrating that PA-Tool consistently achieves the best performance across all models and tasks on both MetaTool and RoTBench benchmarks.
5.1 MetaTool: Consistent Improvements Across All Subtasks
Table 1 shows that PA-Tool consistently improves tool selection performance across all four MetaTool subtasks. In the Similar and Scenario tasks, PA-Tool achieves improvements of up to 10.7 % points, demonstrating that even simple alignment of tool names with the model’s internal knowledge effectively enhances tool selection. This phenomenon is more pronounced in the Multi-tool task, where gains reach up to 9.6 % points, with Llama3.1-8B improving from 78.7% to 88.3%. When tasks require identifying multiple tools simultaneously, the impact of schema misalignment compounds across each tool selection, making alignment particularly critical.
The Reliability task shows substantial improvements of up to 17.0%. For example, Llama3.2-3B improves from 43.6% to 60.6%. This task requires models to recognize when no suitable tool exists in the candidate list. Success depends on accurately understanding what each available tool does. PA-Tool aligns tool names with pretrained knowledge, enabling models to more clearly distinguish between available options and confidently determine when none match the query requirements.
Notably, the Greedy Schema baseline occasionally underperforms the Base model, with Llama3.2-3B showing 39.8% versus 43.6% in Reliability. This can be attributed to greedy decoding generating only a single schema candidate, which limits exploration of the schema space and increases susceptibility to suboptimal or misaligned schemas.
5.2 RoTBench: Persistent Benefits in Multi-Turn Interactions
@@
RoTBench evaluates models in both single-turn and multi-turn conversational settings across two sequential stages: tool selection and parameter identification. In tool selection, PA-Tool demonstrates consistent improvements of 5-10% points across all models in single-turn scenarios. Importantly, these benefits remain robust in multi-turn settings with gains of 4-10% points in multi-turn settings, demonstrating that schema alignment remains effective across extended conversational contexts.
Parameter identification also shows improvements, though more modest than tool selection. The most notable gain appears in multi-turn settings, where Qwen2.5-7B nearly doubles from 5.7% to 11.4%. The smaller overall improvement stems from the nature of parameter naming. Parameters typically have explicit, descriptive names that already align well with common conventions (e.g., user_id with description "a unique user identifier"). Unlike tool names, which vary significantly across different tools, parameter names follow more standardized patterns. This leaves less room for alignment-based improvements.
5.3 Comparison with Closed-source Models
While our open-source models with PA-Tool show substantial improvements, a performance gap remains compared to state-of-the-art closed-source models. Claude Sonnet 4.5 achieves 83.3% tool selection and 67.6% parameter identification on RoTBench single-turn, significantly outperforming even our best configurations. However, PA-Tool narrows this gap, elevating SLMs closer to closed-source baselines while maintaining the computational efficiency advantages of SLMs.
@@ 5.3 Comparison with Closed-source Models
Notably, in certain aspects of tool selection, PA-Tool enables small models to achieve competitive or superior performance. In MetaTool’s Multi-tool task, Llama3.1-8B with PA-Tool achieves 88.3%, surpassing all closed-source models including Claude Sonnet 4.5. Similarly, in the Reliability task, Qwen2.5-7B with PA-Tool reaches 88.2%, closely approaching Gemini 2.5 Flash’s 89.2%. These results demonstrate that schema alignment can enable resource-efficient models to match or exceed state-of-the-art systems in specific aspects of tool selection, particularly in tasks where schema misalignment is a primary failure mode.
@@ 6.1 Error Analysis
Figure 3: Error type distribution for Llama3.1-8B on MetaTool tool selection tasks.
To understand how PA-Tool addresses different failure modes, we analyze error type distributions in Llama3.1-8B before and after applying PA-Tool. We categorize all incorrect predictions from both Base and PA-Tool configurations into three error types: Schema Misalignment Error, where the model generates non-existent or incorrectly named tools; Functional Confusion Error, where the model selects wrong tools with similar functionality; and Context Understanding Error, where the model selects functionally unrelated tools. Detailed error type definitions and experimental settings are provided in Appendix E.
Results. Figure 3 shows how PA-Tool affects different error types. Schema Misalignment Errors decrease dramatically by 80.0% (from 80 to 16 cases), confirming that aligning tool names with pretrained knowledge directly addresses this primary failure mode. In contrast, Functional Confusion Errors decrease by 24.0% (from 271 to 206 cases) and Context Understanding Errors decrease by 18.8% (from 607 to 493 cases). This stark difference—80.0% reduction for schema-related errors versus 18.8-24.0% for other types—demonstrates PA-Tool’s targeted impact. Schema alignment effectively eliminates naming-related failures, but distinguishing functionally similar tools and understanding complex query contexts require capabilities beyond schema optimization.
@@ 6.2 Integration with Supervised Fine-tuning
While PA-Tool is training-free, we demonstrate it remains effective when combined with supervised fine-tuning (SFT), addressing a key question: do models retain pretraining-aligned naming preferences after task-specific adaptation? Detailed information about SFT is presented in Appendix D. We compare four configurations: (1) Vanilla: base model with original schemas; (2) SFT only: fine-tuned model with original schemas; (3) PA-Tool only: base model with aligned schemas; (4) SFT + PA-Tool: SFT model with schemas aligned using the SFT model’s own PA-Tool generation.
Results. Table 2 shows PA-Tool provides consistent gains whether applied alone or with SFT. Notably, PA-Tool alone (third row) outperforms SFT alone (second row) across Similar, Reliability and Multi-tool by up to 10.1%, demonstrating superior efficiency without training. When combined, SFT + PA-Tool (fourth row) achieves the best performance on Similar (72.7%), Scenario (80.8%) and Multi-tool (89.9%) with consistent improvements over SFT alone across all subtasks.
Implications. The consistent gains demonstrate that pretraining-aligned naming preferences persist through fine-tuning. Schemas generated by PA-Tool using either the base or fine-tuned model are highly similar and both improve the model’s performance, indicating that task-specific adaptation does not override deeply internalized naming conventions from pretraining. This shows PA-Tool can deliver further performance gains on top of fine-tuning: schema-level alignment remains necessary to bridge the gap between arbitrary documentation standards and the model’s pretrained knowledge.
Configuration Similar Scenario Reliability Multi-tool
Llama3.1-8B 65.2 71.1 55.1 79.8
+ SFT 71.2 77.6 57.1 82.8
+ PA-Tool 72.2 75.9 67.2 89.9
+ SFT + PA-Tool 72.7 80.8 57.1 89.9
@@
Table 2: Results of fine-tuning method and PA-Tool on MetaTool 10% random sampled test set.
6.3 Hyperparameter Analysis
We investigate the impact of PA-Tool’s three key hyperparameters on performance: the number of candidates ($N$), the similarity threshold ($\alpha$), and the sampling temperature ($t$). All experiments are conducted on the MetaTool benchmark, and we report the average accuracy across all four subtasks. Detailed results are provided in Appendix F.
Number of candidates ($N$) (Figure 4, Top). Smaller models (3B) achieve stable performance with 16-32 candidates, while larger models (7-8B) require 32-64 candidates before performance plateaus. Beyond these ranges, additional candidates provide minimal gains.
Similarity threshold ($\alpha$) (Figure 4, Middle). Performance peaks at $\alpha=0.2$ consistently across all models. At $\alpha=0.1$, performance drops by 2-3 % points, while $\alpha=0.3$ shows similar degradation.
Sampling temperature ($t$) (Figure 4, Bottom). Performance remains stable across temperatures $t\in[0.2,1.0]$, varying within 1-2 % points. Moderate temperatures (t=0.4t=0.4-$0.6$) show slightly better results, though the differences are marginal.
@@ 7 Conclusion
We introduced PA-Tool, a training-free method that aligns tool schemas with models’ pretrained knowledge by leveraging peakedness as a pretraining familiarity signal. Experiments demonstrate improvements of up to 17% points with schema misalignment errors reduced by 80%, validating schema adaptation as an effective strategy for enhancing tool use in small language models. PA-Tool’s practical advantages make it particularly valuable for resource-constrained deployments—as a simple schema-level intervention, it can be applied without model training or fine-tuning, requiring only straightforward name mapping. By bridging pretrained knowledge and real-world tool interfaces, PA-Tool unlocks small models’ potential for tool-augmented applications while preserving computational efficiency.
@@
While PA-Tool effectively reduces schema misalignment errors by 80%, it shows more modest improvements for functional confusion and context understanding errors, indicating that schema optimization alone cannot address all failure modes. Our reliance on peakedness as an alignment signal assumes this metric reliably indicates pretraining familiarity. While validated across our experiments, this relationship may vary for models with substantially different training distributions. Additionally, our evaluation focuses on English-language schemas; the effectiveness of character-level similarity metrics may differ for non-Latin scripts or morphologically complex languages. PA-Tool incurs one-time computational overhead from candidate generation and evaluation. While negligible compared to fine-tuning, this cost could become significant for tool libraries with thousands of components.

### v2

@@ 5 Main Results
PA-Tool substantially improves performance over Base models across all four MetaTool subtasks. The most substantial gains appear in Reliability (up to 17.0%, e.g., Llama3.2-3B: 43.6→60.6%), where models must recognize when no suitable tool exists, a task critically dependent on clearly understanding available options. In Multi-tool, gains reach 9.6% (Llama3.1-8B: 78.7→88.3%). When tasks require identifying multiple tools simultaneously, schema misalignment compounds across each selection, making alignment particularly critical. Similar and Scenario tasks show improvements up to 10.7%.
PA-Tool consistently outperforms training-free alternatives. Greedy occasionally underperforms Base (Llama3.2-3B Reliability: 39.8% vs. 43.6%) by generating only a single candidate, limiting schema space exploration. While MostFreq sometimes achieves competitive results by capturing frequency, PA-Tool shows more consistent improvements by measuring distributional concentration. Compared to Human-designed schemas, PA-Tool achieves comparable or superior performance (Llama3.2-3B Similar: 65.7% vs. 58.6%), demonstrating automated alignment can match human intuition while being scalable.
@@
RoTBench evaluates models in single-turn and multi-turn settings across tool selection and parameter identification. PA-Tool demonstrates consistent improvements of 5-10% in single-turn tool selection, with gains of up to 10% remaining in multi-turn settings, demonstrating effectiveness across extended conversations.
MostFreq shows strong performance for some models in single-turn (Qwen2.5-3B: 12.4%), but PA-Tool demonstrates more consistent improvements in multi-turn settings and across larger models, suggesting better generalization.
Comparison with Closed-source Models.
@@
While a performance gap remains versus state-of-the-art closed-source models (Claude Sonnet 4.5: 83.3% tool selection on RoTBench single-turn), PA-Tool enables small models to achieve competitive or superior performance in specific aspects. In MetaTool’s Multi-tool, Llama3.1-8B with PA-Tool (88.3%) surpasses all closed-source models including Claude Sonnet 4.5 (85.1%). In Reliability, Qwen2.5-7B with PA-Tool (88.2%) approaches Gemini 2.5 Flash (89.2%). These results demonstrate that targeted schema alignment can enable resource-efficient models to match state-of-the-art systems where naming conventions critically impact performance.
Generalization to Diverse Models.
Beyond the models examined here, we conduct additional experiments on diverse model families and scales, including both SLMs (Ministral-8B, GPT-4.1-nano) and LLMs (Llama-3.3-70B, GPT-4.1-mini, Gemini-2.5-Flash). PA-Tool shows substantial improvements for SLMs with consistent gains up to 11.2% on RoTBench. For LLMs, improvements are generally more modest, though notable exceptions exist—GPT-4.1-mini achieves 12.0% gain on MetaTool Multi-tool, demonstrating that even highly capable models benefit from schema alignment in certain tasks (see Appendix E for details).
@@ 6.1 Error Analysis
Figure 3: Error type distribution for Llama-3.1-8B on MetaTool tool selection tasks.
To understand how PA-Tool addresses different failure modes, we analyze error distributions in Llama3.1-8B before and after applying PA-Tool. We categorize incorrect predictions into three types: Schema Misalignment Error (generating nonexistent but plausible tools), Functional Confusion Error (selecting wrong tools with similar functionality), and Context Understanding Error (selecting functionally unrelated tools). Detailed definitions and settings are in Appendix G.
Figure 3 shows how PA-Tool affects different error types. PA-Tool substantially reduces all error types, with particularly strong impact on Schema Misalignment (80.0% reduction) and meaningful improvements on Functional Confusion (24.0%) and Context Understanding (18.8%). This demonstrates that schema alignment not only eliminates naming-related failures but also indirectly improves the model’s ability to distinguish tools and understand contexts, likely by reducing cognitive load from misaligned names.
6.2 Integration with Training-Free Methods
Beyond schema renaming, we evaluate PA-Tool’s synergy with other training-free approaches. We test three categories of methods: (1) Retrieval-based correction, which maps misaligned outputs to valid tools post-hoc using BM25 (Robertson and Zaragoza, 2009) or ToolLLM (Qin et al., 2024); (2) Constrained generation, which enforces valid tool names through JSON schema constraints during decoding; and (3) Description enhancement (EasyTool (Yuan et al., 2025)), which rewrites tool descriptions for clarity without modifying names. These methods address different aspects of tool-use and are compatible with PA-Tool’s schema alignment approach. We evaluate both standalone and combined configurations (Method + PA-Tool), with detailed settings presented in Appendix K.
Table 11 shows that retrieval methods show limited gains on their own (<3% improvement). Even with post-hoc correction opportunities, matching often fails to capture the model’s intended tool. Constrained generation substantially improves RoTBench (Qwen2.5-7B: 78.1% vs. Base 49.5%) by eliminating format errors, but shows mixed results on MetaTool.
However, integrating PA-Tool with these methods consistently improves both settings (Table 12), confirming the two approaches are complementary. PA-Tool consistently outperforms EasyTool (Table 13), with notable gaps on MetaTool (e.g., Llama3.2-3B Similar: 65.7% vs. 45.3%). This suggests naming alignment is more effective than description enhancement. Moreover, rewriting descriptions risks altering perceived tool functionality, while PA-Tool preserves semantics by modifying only names. Nevertheless, both remain compatible, with some combined configurations showing competitive results. Across all methods, PA-Tool demonstrates broad compatibility and provides complementary gains when combined, confirming it addresses a distinct failure mode.
@@ 6.3 Integration with Supervised Fine-tuning
[…]
Table 2 shows PA-Tool provides consistent gains across all configurations. Notably, PA-Tool alone outperforms both SFT1 and SFT2 on Reliability and Multi-tool without any training. SFT2 shows marginal gains over SFT1 on MetaTool but degrades on RoTBench (59.1% vs. 61.0%), indicating catastrophic forgetting. Combined configurations achieve the best results, with SFT2 + PA-Tool reaching 73.2% on Similar. These results reveal two key insights. First, pretraining-aligned naming preferences persist through fine-tuning, as schemas generated from base or fine-tuned models are similarly effective. Second, SFT’s cross-domain degradation highlights its limitation: it specializes to specific tool distributions, risking catastrophic forgetting. PA-Tool’s training-free nature avoids this, providing robust improvements across both in-domain and out-of-domain benchmarks.
Configuration MetaTool RoTBench
Similar Scenario Reliab. Multi. Tool Sel.
Base 65.2 71.1 55.1 79.8 58.1
+ PA-Tool 72.7 75.9 67.2 89.9 68.6
+ SFT1 71.2 77.6 57.1 82.8 61.0
+ SFT1 + PA-Tool 72.7 80.8 57.1 89.9 65.7
[…]
+ SFT2 + PA-Tool 73.2 79.2 55.6 86.9 61.9
@@
Table 2: Results of fine-tuning and PA-Tool on 20% random sample from MetaTool and full RoTBench.
Figure 4: Peakedness across training epochs.
6.4 Validating Peakedness
PA-Tool assumes that peakedness reflects models’ familiarity with naming patterns. To validate this, we simulate pretraining conditions where models encounter tool schemas and measure whether peakedness increases across training epochs. We use the same MetaTool training data and experimental settings from Section 6.3 but with a different loss computation: unlike SFT which computes loss only on assistant responses, we compute loss on all tokens to simulate pretraining conditions. We train four models (Qwen2.5-3B/7B, Llama3.2-3B, Llama3.1-8B), measuring peakedness at epochs 0, 10, 20, 30, 50, and 100 with 256 candidates per tool component for stable measurements. Figure 4 shows that peakedness consistently increases with training epochs across all models, with gains up to +25.8% (Llama3.2-3B). This validates our hypothesis that peakedness captures learning exposure, supporting PA-Tool’s use of peakedness to identify pretraining-aligned naming conventions.
6.5 Additional Benchmarks
To demonstrate that PA-Tool generalizes across different evaluation settings, we evaluate on API-Bank Li et al. (2023b) and $\tau$-Bench Yao et al. (2025). API-Bank evaluates tool-calling accuracy by measuring exact matches between predicted and ground-truth tools and parameters. $\tau$-Bench is an end-to-end benchmark that assesses agent performance through multi-turn dialogues where agents must complete tasks through interaction with users.
API-Bank. PA-Tool consistently improves performance in most settings (Table 3). Qwen2.5-3B shows substantial gains on the Call task (18.0→28.5), while improvements persist on Call+Retrieve, which requires both tool selection and information retrieval. This confirms that schema alignment enhances tool-calling accuracy.
$\tau$-Bench. To verify that improved tool selection translates to better task completion, we evaluate on $\tau$-Bench (Retail). We run each evaluation 5 times to mitigate variance from the user simulator. PA-Tool consistently improves end-to-end task completion across all models (Table 3). Analyzing Llama3.1-8B’s tool-calling trajectories reveals that schema misalignment errors decreased from 115 to 98 cases, demonstrating that addressing schema misalignment contributes substantially to the observed performance gains.
Call Call+Retrieve $\tau$-Bench
Model Base PA-Tool Base PA-Tool Base PA-Tool
Qwen2.5-3B 18.0 28.5 18.5 21.9 3.7 3.8
Qwen2.5-7B 25.7 34.7 23.5 26.9 6.8 9.7
Llama3.2-3B 5.1 5.4 10.9 9.2 4.5 5.6
Llama3.1-8B 28.0 29.8 21.0 22.4 9.7 11.1
Table 3: Performance on API-Bank and end-to-end task completion (%) on $\tau$-Bench (Retail, N=5).
@@ 7 Conclusion
We introduced PA-Tool, a training-free method that aligns tool schemas with models’ pretrained knowledge by leveraging peakedness as a pretraining familiarity signal. Experiments demonstrate improvements of up to 17% with schema misalignment errors reduced by 80%, validating schema adaptation as an effective strategy for enhancing tool use in small language models. PA-Tool’s practical advantages make it particularly valuable for resource-constrained deployments—as a simple schema-level intervention, it can be applied without model training or fine-tuning, requiring only straightforward name mapping. By bridging pretrained knowledge and real-world tool interfaces, PA-Tool unlocks small models’ potential for tool-augmented applications while preserving computational efficiency.
@@
The effectiveness of PA-Tool varies with model capacity in an interpretable way: smaller models show consistent improvements, as schema misalignment remains a primary failure mode. In contrast, larger models exhibit diminishing gains as parameter count increases, likely due to their stronger reasoning abilities. However, performance improvements remain model-dependent; GPT-4.1-mini still achieves a 12.0% improvement on the Multi-tool task in Metatool, suggesting that certain architectures benefit from schema alignment regardless of scale.
Our reliance on peakedness as an alignment signal assumes this metric reliably indicates pretraining familiarity. While validated across our experiments, this relationship may vary for models with substantially different training distributions. Additionally, our evaluation focuses on English-language schemas; the effectiveness of character-level similarity metrics may differ for non-Latin scripts or morphologically complex languages. PA-Tool incurs one-time computational overhead from candidate generation and evaluation. While negligible compared to fine-tuning, this cost could become non-trivial for tool libraries with thousands of components.
@@ Appendix B Prompt Templates
Training Data Generation Prompt You are an expert at explaining tool selection reasoning for AI systems. Your task is to generate high-quality, flowing reasoning that explains WHY specific tools were chosen for a given query. Your Task: Given: A user’s query A list of available tools with descriptions The tools that were selected (answer) Generate: A clear, logical reasoning in natural paragraph form that explains why these specific tools are the best choices. Quality Criteria: Your reasoning MUST: 1. Analyze the query deeply: Identify key requirements, implicit needs, and user intent 2. Connect query to tools: Explicitly link query elements to tool capabilities 3. Justify each selection: Explain why each tool is necessary 4. Show tool synergy: Explain how the tools work together 5. Consider alternatives: Briefly mention why other tools were NOT chosen 6. Be specific: Use concrete details from the query and tool descriptions 7. Write in flowing prose: Use natural paragraphs, not bullet points or lists Example: Now Generate: User’s Query: {{ query }} Selected Tools: {{ selected_tools }} Write your reasoning as flowing paragraphs that naturally cover query analysis, tool justification for each selected tool, their combined strategy, and why alternatives were not chosen.
Figure 7: Training data generation prompt for supervised fine-tuning in Metatool.
Prompt template used for inference on MetaTool You are a helpful AI assistant. Your current task is to choose the appropriate tool to solve the user’s query based on their question. I will provide you with the user’s question and information about the tools. If there is a tool in the list that is applicable to this query, please return the name of the tool (you can only choose one tool). If there isn’t, please return ‘None.’ Additionally, you will need to support your answer with a brief explanation. User’s Query: [User’s Query Start] Planning a beach day next week, while considering the ideal weather conditions, availability of beach accessories, preferred beach location, potential activities to engage in like swimming, sunbathing, or playing beach games, and coordinating with friends or family members to join in the fun. [User’s Query End] List of Tools with Names and Descriptions: [List of Tools with Names and Descriptions Start] 1. tool name: airqualityforeast, tool description: [‘Planning something outdoors? Get the 2-day air quality forecast for any US zip code.’] 2. tool name: WeatherTool, tool description: Provide you with the latest weather information. 3. tool name: AusSurfReport, tool description: [‘Get today’s surf report for any break throughout Australia!’] 10. tool name: EarthquakeTool, tool description: Provides real-time earthquake notifications and news. [List of Tools with Names and Descriptions End] Here are some examples: [Examples Start] query: “I’m planning a hiking trip next week. What will the weather be like in the Grand Canyon?” tool: WeatherTool query: “I’m a beginner surfer. Can you suggest a beach with mild waves for me to practice in Australia?” tool: AusSurfReport [Examples End] User query: “Planning a beach day next week, while considering the ideal weather conditions…” tool:
Figure 8: Prompt template used for model inference on MetaTool benchmark. The prompt includes task instructions, a list of candidate tools with descriptions, few-shot examples, and the target user query.
Prompt template used for model inference on RoTBench System: You are an expert in using tools. At each step, analyze the state and decide the next action with a function call. Available tools: [{"name": "get_translation_nllb", "description": "Translate text using NLLB model.", "parameters": {"input_text": {...}, "tgt_lang": {...}, "src_lang": {...}, ...}}, {"name": "get_translation_baidu", "description": "Translate using BAIDU API.", "parameters": {"text": {...}, "tgt_lang": {...}, ...}}, ...] User: “What is the translation of ‘See you later’ in Japanese?” Expected output: Thought: Need to translate English to Japanese Action: get_translation_nllb Action Input: {"input_text": "See you later", "tgt_lang": "jpn_Jpan", "src_lang": "eng_Latn", …}
Figure 9: Prompt template used for model inference on RoTBench. The prompt includes a system message, available tool schemas with full parameter specifications, and the user query. Models must output structured responses with reasoning, tool selection, and parameter values.
Prompt template used for model inference on API-Bank Instruction: Based on the given API description and the existing conversation history 1..t, please generate the API request that the AI should call in step t+1 and output it in the format of [ApiName(key1=’value1’, key2=’value2’, …)], replace the ApiName with the actual API name, and replace the key and value with the actual parameters. Your output should start with a square bracket “[” and end with a square bracket “]”. Do not output any other explanation or prompt or the result of the API call in your output. This year is 2023. API descriptions: [{"name": "GetWeather", "description": "Get current weather information", "parameters": {"location": {"type": "string", "description": "City name"}, "unit": {"type": "string", "enum": ["celsius", "fahrenheit"]}}}, ...] Input: User: What’s the weather like in San Francisco? AI: I’$l$ check the current weather in San Francisco for you. Expected output: [GetWeather(location=’San Francisco’, unit=’fahrenheit’)]
Figure 10: Prompt template used for model inference on API-Bank. The prompt includes instructions for API call generation based on conversation history, API schemas with parameter specifications, and example dialogue with expected API call output in bracket notation.
Prompt template used for model inference on Tau-Bench Instruction: You need to act as an agent that uses the above tools to help the user according to the above policy. At each step, your generation should have exactly the following format: Thought: <A single line of reasoning to process the context and inform the decision making. Do not include extra lines.> Action: {"name": <The name of the action>, "arguments": <The arguments to the action in json format>} The Action will be parsed, so it must be valid JSON. You should not use made-up or placeholder arguments. Available tools: [{"type": "function", "function": {"name": "get_current_weather", "description": "Get the current weather", "parameters": {"type": "object", "properties": {"location": {"type": "string", "description": "The city and state, e.g. San Francisco, CA"}, "format": {"type": "string", "enum": ["celsius", "fahrenheit"], "description": "The temperature unit to use. Infer this from the users location."}}, "required": ["location", "format"]}}}, ...] User: “I want to know the current weather of San Francisco” Expected output: Thought: Since the user asks for the weather of San Francisco in USA, the unit should be in fahrenheit. I can query get_current_weather to get the weather. Action: {"name": "get_current_weather", "arguments": {"location": "San Francisco, CA", "format": "fahrenheit"}} [If tool returns "70F"] Thought: I can answer the user now. Action: {"name": "respond", "arguments": {"response": "The current weather of San Francisco is 70F."}}
Figure 11: Prompt template used for model inference on Tau-Bench. The prompt includes instructions for structured agent behavior, available tool schemas with parameter specifications, and example user query with expected multi-step reasoning and tool invocations.